OpenStudy (anonymous):

Complete the equation of the graphed linear function in point-slope form. y – 8 =[ ](x –[ ])

OpenStudy (anonymous):

Last question @iGreen.

OpenStudy (anonymous):

\(y - y_1 = m(x - x_1)\) Where y1 is the y-value of the point, x1 is the x-value of the point, and 'm' is the slope.

OpenStudy (anonymous):

It gives you the point (2, 8). So x1 here is 2, and y1 here is 8. Can you plug that into point-slope form? \(y - y_1 = m(x - x_1)\)

OpenStudy (anonymous):

Can you do it?

OpenStudy (anonymous):

y= 2 x=6?

OpenStudy (anonymous):

x1 = 2, y1 = 8

OpenStudy (anonymous):

oh woops XP

OpenStudy (anonymous):

Try again.

OpenStudy (anonymous):

(6,0)

OpenStudy (anonymous):

It gives you the point (2, 8). So x1 here is 2, and y1 here is 8. \(y - y_1 = m(x - x_1)\) \(y - 8 = m(x - 2)\) Now let's find the slope..

OpenStudy (anonymous):

Take any two points: (6, 0) (2, 8) Plug them in: \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{8-0}{2-6}\) Simplify

OpenStudy (anonymous):

(8,-4)

OpenStudy (anonymous):

Yes, so the slope is -8/4, simplify it.

OpenStudy (anonymous):

Just divide -8 / 4

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

Yes, so our final answer is \(y - 8 = -2(x - 2)\)

OpenStudy (anonymous):

thank you again :P

OpenStudy (anonymous):

No problem :P